OpenStudy (anonymous):

if a and 1/a are the zeroes of the poly. 4x^-2x+(k-4),find the value of k.

OpenStudy (lgbasallote):

x = a and x = 1/a so the factors are (x-a)(ax - 1) let's FOIL ax^2 - x - a^2x + a ax^2 - x - a^2x + a = 4x^2 - 2x+ k - 4 ax^2 = 4x^2 therefore a = 4 a = k - 4 4 = k - 4 therefore k = 0 that right?

OpenStudy (anonymous):

ya